Solve the system by the elimination method.
5x + 3y = -11
7x - 2y = 19

Answer by mananth(16949)

5 x + 3 y = -11 .............1
7 x -2 y = 19 .............2
Eliminate y
multiply (1)by 2
Multiply (2) by 3
10 x 6 y = -22
21 x -6 y = 57
Add the two equations
31 x = 35
/ 31
x = 1.13
plug value of x in (1)
5 x + 3 y = -11
5.65 + 3 y = -11
3 y = -11 -5.65
3 y = -16.65
y = -5.55